OpenStudy (anonymous):

Teesha is in the French club. There are 29 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will not be picked as a guide? (1 point)

OpenStudy (anonymous):

picked first - P = 1/31 picked second - P = 30/31 * 1/30 = 30/930 = 1 / 31 required P = 1/31 + 1/31 = 2/31

OpenStudy (anonymous):

2/31 the answer